# Principal stratification with recurrent events truncated by a terminal event: A nested Bayesian nonparametric approach

### v2

@@ Appendix D Details of Gibbs sampler
For simplicity, in what follows, we denote by $\mathbf{X}_{i}$ the augmented covariates that contain the baseline covariates and treatment variable $Z_{i}$ and denote by $\mathbf{X}_{ij}$ the covariates that additionally include the time-varying covariates at the $j$-th recurrent period $\mathbf{V}_{ij}$.
D.1 Imputation of $U$
@@
$U_{i}\mid-\sim\mathrm{TN}\quantity(\mathbf{X}_{i}^{\top}\beta_{u,G_{i}}+\gamma^{Z_{i}}_{G_{i}},\tau_{G_{i}}^{2},U_{i}^{\mathrm{obs}},\infty)),$
@@ D.2 Imputation of $Y_{i(N_{i}+1)}$
D.3 Update of $\mathbf{G}$ and $\mathbf{H}$
@@
$p\quantity(H_{ij}=l\mid-)\propto w_{l\mid G_{i}}^{\theta}\times\phi\quantity(Y_{ij}\Big|\mathbf{X}_{ij}\beta_{y,l\mid G_{i}}+\psi_{l\mid G_{i}}\gamma^{Z_{i}}_{G_{i}},\sigma_{l\mid G_{i}}^{2}),$
@@ D.4 Update of $w^{\phi}_{k}$ and $v^{\phi}_{k}$
$v^{\phi}_{k}\sim\text{Be}\left(1+\sum_{i=1}^{n}\mathbbm{1}(G_{i}=k),\alpha+\sum_{i=1}^{n}\mathbbm{1}(G_{i}>k)\right).$ (16)
@@
D.5 Update of $w^{\theta}_{l\mid k}$ and $v^{\theta}_{l\mid k}$
@@
Then update $w^{\theta}_{l\mid k}=u^{\theta}_{l\mid k}\prod_{j=1}^{l-1}(1-u^{\theta}_{j\mid k})$ for $k=1,\ldots,K$.
D.6 Update of $\alpha_{\phi}$ and $\alpha_{\theta\mid k}$
@@
$\alpha_{\phi}$ $\sim\mathrm{Ga}\left(a_{\alpha}+K-1,b_{\alpha}-\sum_{k=1}^{K-1}\ln(1-w^{\phi}_{k})\right),$
$\alpha_{\theta\mid k}$ $\sim\mathrm{Ga}\left(a_{\alpha}+L-1,b_{\alpha}-\sum_{l=1}^{L-1}\ln(1-w^{\theta}_{l\mid k})\right).$
D.7 Update of $(\beta_{u,k},\tau_{k})$
@@
D.8 Update of $(\beta_{y,l\mid k},\sigma_{l\mid k})$
@@
D.9 Update of $\boldsymbol{\gamma}_{k}$
@@ D.10 Update of $\psi_{l\mid k}$
D.11 Compute estimands
Given all model parameters, we compute $\kappa_{t,r}(z,x,\boldsymbol{\gamma})$, $\zeta_{t,r}(z,x,\boldsymbol{\gamma})$ and $\eta_{r}(z,x,\boldsymbol{\gamma})$. They are not easy to compute in closed form due to the nonparametric nature of the proposed prior. Therefore we approximate them by the Monte carlo simulation within each MCMC iteration. For example, we approximate $\kappa_{t,r}(z,x,\boldsymbol{\gamma})\approx 1/R\sum_{r=1}^{R}N_{i}(t)$ with $R$ simulated samples of $N_{i}(t)$ from its posterior predictive distribution.
Appendix E Sensitivity analysis
Figure 5: Posterior distributions of SANR under five sensitivity parameters ($\rho=0.1,0.3,0.5,0.7,0.9$), shown from top to bottom.
Figure 6: Posterior distributions of SAER under five sensitivity parameters ($\rho=0.1,0.3,0.5,0.7,0.9$), shown from top to bottom.
This section complements the empirical results in Section 5 with additional sensitivity analyses for the unidentifiable frailty parameter. Figures 5 and 6 report posterior surfaces for SANR and SAER obtained under five sensitivity parameters $\rho\in\{0.1,0.3,0.5,0.7,0.9\}$. Overall, varying the sensitivity parameters does not cause substantial shifts in the posterior estimates, so our substantive conclusions remain unchanged.

### v3

@@ Appendix C Details of Gibbs sampler
For simplicity, in what follows, we denote by $\mathbf{X}_{i}$ the augmented covariates that contain the baseline covariates and treatment variable $Z_{i}$ and denote by $\mathbf{X}_{ij}$ the covariates that additionally include the time-indexed covariates at the $j$-th recurrent period $\mathbf{V}_{ij}$.
C.1 Imputation of $U$
@@
$U_{i}\mid-\sim\mathrm{TN}\quantity(\mathbf{X}_{i}^{\top}\beta_{u,G_{i}}+\gamma^{Z_{i}}_{G_{i}},\tau_{G_{i}}^{2},U_{i}^{\mathrm{obs}},\infty),$
@@ C.2 Imputation of $Y_{i(N_{i}+1)}$
C.3 Update of $\mathbf{G}$ and $\mathbf{H}$
@@
$p\quantity(H_{ij}=l\mid-)\propto w_{l\mid G_{i}}^{\theta}\times\mathrm{N}\quantity(Y_{ij}\Big|\mathbf{X}_{ij}\beta_{y,l\mid G_{i}}+\psi_{l\mid G_{i}}\gamma^{Z_{i}}_{G_{i}},\sigma_{l\mid G_{i}}^{2}),$
@@ C.4 Update of $w^{\phi}_{k}$ and $v^{\phi}_{k}$
$v^{\phi}_{k}\sim\text{Be}\left(1+\sum_{i=1}^{n}\mathbbm{1}(G_{i}=k),\alpha_{\phi}+\sum_{i=1}^{n}\mathbbm{1}(G_{i}>k)\right).$ (14)
@@
C.5 Update of $w^{\theta}_{l\mid k}$ and $v^{\theta}_{l\mid k}$
@@
Then update $w^{\theta}_{l\mid k}=v^{\theta}_{l\mid k}\prod_{j=1}^{l-1}(1-v^{\theta}_{j\mid k})$ for $k=1,\ldots,K$.
C.6 Update of $\alpha_{\phi}$ and $\alpha_{\theta\mid k}$
@@
$\alpha_{\phi}$ $\sim\mathrm{Ga}\left(a_{\alpha}+K-1,b_{\alpha}-\sum_{k=1}^{K-1}\ln(1-v^{\phi}_{k})\right),$
$\alpha_{\theta\mid k}$ $\sim\mathrm{Ga}\left(a_{\alpha}+L-1,b_{\alpha}-\sum_{l=1}^{L-1}\ln(1-v^{\theta}_{l\mid k})\right).$
C.7 Update of $(\beta_{u,k},\tau_{k})$
@@
C.8 Update of $(\beta_{y,l\mid k},\sigma_{l\mid k})$
@@
C.9 Update of $\boldsymbol{\gamma}_{k}$
@@ C.10 Update of $\psi_{l\mid k}$
C.11 Compute estimands
Given all model parameters, we compute $\kappa_{t,r}(z,x,\boldsymbol{\gamma})$ and $\eta_{r}(z,x,\boldsymbol{\gamma})$. They are not easy to compute in closed form due to the nonparametric nature of the proposed prior. Therefore we approximate them by the Monte Carlo simulation within each MCMC iteration. For example, we approximate $\kappa_{t,r}(z,x,\boldsymbol{\gamma})\approx 1/B\sum_{b=1}^{B}N_{i}^{(b)}(t)$ with $B$ simulated samples of $N_{i}^{(b)}(t)$ from its posterior predictive distribution.
Appendix D Contour plots for SANR
Figure 5 presents contour plots for the SANR across various correlation parameters $\rho$. These plots illustrate how SANR values evolve over a grid of time points $r$ and $t$ ($r\leq t$), consistent with the sensitivity analysis patterns observed in the main manuscript.
Figure 5: Contour plots of posterior mean (left), $95\%$ posterior lower bound (middle), and upper bound (right) for SANR. The posterior values are computed for different cut-off values of $t$ and $r$ from $360$ days to $1440$ days with an increment of $90$ such that $t\leq r$, and interpolated between the grid values with a linear spline. The regions with the estimate greater than 11 are represented using warm colors (yellow to red), while regions with the estimate less than 11 are indicated using cool colors (blue).